# Temperature-dependent Photoluminescence and Raman Spectroscopy of Selenium Thin Film Solar Cells

### v1

@@ I Introduction
Selenium is a wide-bandgap elemental semiconductor that has attracted renewed interest for a range of optoelectronic applications, including photodetectors [1, 2, 3], radiation sensors [4, 5], and, most notably, indoor [6, 7, 8, 9, 10, 11] and tandem photovoltaics [12, 13, 14, 15, 16]. Its quasi-one-dimensional trigonal crystal structure, shown in Figure 1(a), gives rise to pronounced anisotropy in material properties, enabling more advanced technologies such as polarization-sensitive light sources and detectors [17, 18, 19, 20]. In addition to its low-dimensional structure, trigonal selenium offers a direct bandgap in the 1.8-2.2 eV range [21, 22, 23], which is tunable through alloying with tellurium [24, 25, 26], a high optical absorption coefficient [27, 6], tolerance to intrinsic point defects [28, 29, 30], low-temperature processing [31, 32, 33, 34], long-term air stability [35, 36], and the irresistible simplicity of a single-element system. Yet despite these advantages, the optoelectronic quality of selenium thin films remains a major bottleneck, limiting progress toward high-efficiency photovoltaic devices. Given that photoluminescence (PL) and Raman spectroscopy is widely recognized as the most powerful non-contact characterization tools for evaluating optoelectronic and structural quality [37, 38, 39], a systematic investigation of the radiative emission properties and microstructural quality using these techniques is a natural next step toward identifying both intrinsic and synthesis-induced performance limitations in state-of-the-art selenium thin films.
@@
In contrast to free-standing single crystals, selenium thin films are typically grown directly on substrates and are thus more susceptible to substrate-induced stress, anisotropic grain growth, and structural disorder – all of which can significantly alter PL behavior. In 2022, Nielsen et al. reported two distinct PL emission bands at cryogenic temperatures from a selenium thin film integrated into a record-performing photovoltaic device [27]. Liu et al. observed similar spectral features, although the acquisition conditions were not specified [45]. Other studies have reported redshifted PL peak energies well below the optical bandgap in films infiltrated into mesoporous oxide scaffolds [26, 46]. Li et al. presented PL spectra with more reasonable peak energies, but without a detailed analysis of the nature of the underlying transition [1]. This inconsistency across the literature underscores the need for a more systematic investigation. However, the high vapor pressure of selenium makes thin films prone to sublimation and degradation during PL measurements, even under low excitation conditions. This challenge may have limited the widespread use of PL as a diagnostic tool, despite its value for evaluating optoelectronic quality.
In this work, we use temperature-dependent PL and Raman spectroscopy to study the optoelectronic and vibrational properties of selenium thin films synthesized in different laboratories using state-of-the-art fabrication methodologies. To address the challenge posed by the high vapor pressure and volatility of selenium, we adopt a recently developed closed-space encapsulation strategy introduced by Nielsen et al. [47], which effectively suppresses sublimation during the measurements. The vibrational properties are studied through a detailed analysis of the thermal evolution of Raman peaks, supported by first-principles phonon dispersion calculations that enable identification of all Raman-active modes and their associated atomic displacements. Temperature-dependent PL measurements, complemented by power-dependent excitation studies, are used to examine the nature of two emission bands and quantify the electron-phonon interactions that drive the thermal shifts and broadening of the observed radiative features. By comparing samples produced in different labs, we show that the short-range electron-phonon coupling strength is highly sensitive to subtle processing variations, demonstrating that a reduction of microstructural disorder through careful control of growth parameters and thermal processing can substantially improve the optoelectronic quality of selenium thin films.
@@ II Results
Building on this group-theoretical framework, we next examine the phonon dispersion using first-principles density functional theory (DFT) calculations. Figure 2 shows the phonon dispersion and the projected phonon density of states (DOS), with vibrational modes at the -point labeled by the symmetries introduced above. The A and doubly degenerate E(2) modes appear nearly degenerate in frequency at , suggesting they may be difficult to distinguish experimentally. The spectrum also reveals a pronounced phonon bandgap separating low- and high-frequency optical branches. While such a gap restrict available phonon scattering pathways – potentially suppressing phonon-phonon interactions and enhancing radiative efficiency – it may also channel energy from photoexcited carriers into specific high-energy modes, thereby increasing electron-phonon coupling and promoting non-radiative recombination. To modify or eliminate the phonon gap and thereby tailor intrinsic vibrational properties, phonon dispersion engineering through alloying, nanostructuring, or deliberate introduction of strain offers a promising strategy for enhancing optoelectronic performance potential.
Next, we deconvolute the experimental Raman spectrum acquired at room temperature using a rigorous methodology as detailed in Refs [51, 52], shown in Figure 3(a). The low-frequency peak at 146 cm corresponds well with the doubly degenerate E(1) mode predicted by the phonon dispersion. The peak at 235 cm overlaps with both the predicted E(2) and A modes, which are nearly degenerate and indistinguishable in energy. The slight asymmetry of this peak, particularly evident in the power-dependent Raman spectra in Figure 3(b), suggests contributions from two distinct but overlapping vibrational modes. However, due to the finite linewidth, limited spectral resolution, and low signal intensity, the two modes are best modeled using a single Lorentzian. These assignments are consistent with prior observations by Geick et al. [53]. Additionally, a broad multi-phonon band is observed between 430 and 475 cm, consistent with multi-phonon scattering involving 2A, 2E(2) and A + E(2) modes. The DFT-calculated phonon displacements of the three Raman-active modes are shown as an inset in Figure 3(a), illustrating that the E(1) mode is a bending type vibration that shears the covalent bonds, whereas both E(2) and A correspond to asymmetric and symmetric bond-stretching vibrations, respectively. To better visualize these displacements, additional representations along different lattice directions are provided in Figure S1 of the Supplementary Information.
@@
As the temperature decreases, the E(1) mode is observed to blueshift. This is a vibrational fingerprint of interchain contraction as the restoring force associated with the bond-shearing mode is strengthening, indicating an increased interchain coupling and a reduction in the distance between the selenium chains. In contrast, the E(2)+A peak, arising from two bond-stretching modes, redshifts upon cooling – opposite to what would be expected from simple lattice contraction. Although this redshift is smaller than the instrumental resolution, it is systematic and reproducible. This unexpected softening may result from the relaxation of grown-in compressive strain or the development of tensile strain along the chains at lower temperatures, both of which can reduce the effective restoring forces despite thermal contraction of the Se-Se bonds.
@@ III Discussion
The interplay between vibrational and optoelectronic properties has a strong influence on overall device performance. Our results show that structural disorder in selenium thin films is not intrinsic but highly sensitive to subtle variations in processing. These extended defects likely act as killer non-radiative recombination centers. Because such defects are not necessarily radiative, their impact may not be fully captured by PL measurements alone. However, we do observe a distinct defect-bound exciton emission feature. Prior work by Nielsen et al. [58] identified an acceptor-like trap with a similar energy level, possibly corresponding to the (+1/0) charge transition level of selenium vacancies [28]. While this lower-energy exciton-related emission is unlikely to represent the primary non-radiative loss mechanism responsible for the large open-circuit voltage deficit, it may still contribute to the broad absorption onset commonly observed in both EQE and absorption spectra. This would help explain the persistent discrepancy between our PL-inferred optical bandgaps (1.84–1.86 eV), and the photovoltaic bandgaps (1.95 eV) derived from the inflection point of the EQE [49, 59, 60, 34]. This discrepancy is likely a result of sub-bandgap absorption from band tailing induced by traps and/or structural disorder. Earlier work by Chen et al. [61] also reported strong trapping of free excitons in selenium and proposed that such trapping could account for significant non-radiative recombination losses at higher temperatures. All of these observations are broadly consistent with the emerging picture of the microstructural and optoelectronic properties of selenium thin films.
@@ Acknowledgements
The work presented here is supported by the Carlsberg Foundation, grant CF24-0200. O.S.B. and E.S. acknowledge the CURIO-CITY project (PID2023-148976OB-C41); O.S.B., A.T. and M.P. acknowledge the INNO-PV project (PID2022-140226OB-C31C32); M.P. acknowledges the ENPOWER project (PCI2024-155100-2) from the CETP-Partnership Program 2023; all of them funded by MCIN/AEI/10.13039/501100011033/ FEDER. A.T. and M.P. acknowledge the SELECTRON project (CNS2023-14817) funded by MCIN/AEI/10.13039/501100011033/ NextGenerationEU/PRTR. Through our membership of the UK’s HEC Materials ChemistryConsortium, which is funded by the UK Engineering and Physical Sciences Research Council (EPSRC; EP/L000202, EP/R029431, EP/T022213), this work used the ARCHER2 UK National Supercomputing Services. We are also grateful to the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1).

### v2

# Spectroscopic Signatures of Structural Disorder and Electron-Phonon Interactions in Trigonal Selenium Thin Films for Solar Energy Harvesting

@@ I Introduction
Selenium is a wide-bandgap elemental semiconductor that has attracted renewed interest for a range of optoelectronic applications, including photodetectors [1, 2, 3], radiation sensors [4, 5], and, most notably, indoor [6, 7, 8, 9, 10, 11] and tandem photovoltaics [12, 13, 14, 15, 16]. Its quasi-one-dimensional trigonal crystal structure, shown in Figure 1(a), gives rise to pronounced anisotropy in material properties, enabling more advanced technologies such as polarization-sensitive light sources and detectors [17, 18, 19, 20]. In addition to its low-dimensional structure, trigonal selenium offers a direct bandgap in the 1.8-2.2 eV range [21, 22, 23], which is tunable through alloying with tellurium [24, 25, 26], a high optical absorption coefficient [27, 6], tolerance to intrinsic point defects [28, 29, 30], low-temperature processing [31, 32, 33, 34], long-term air stability [35, 36], and the irresistible simplicity of a single-element system. Yet despite these advantages, the optoelectronic quality of selenium thin films remains a major bottleneck, limiting progress toward high-efficiency photovoltaic devices [27, 28]. Photoluminescence (PL) and Raman spectroscopy is widely recognized as a powerful non-contact characterization tools for evaluating optoelectronic and structural quality of materials [37, 38, 39]. Therefore, a systematic investigation of the radiative emission properties and microstructural quality using these techniques is a natural next step toward identifying both intrinsic and synthesis-induced performance limitations in state-of-the-art selenium thin films.
@@
In contrast to free-standing single crystals, selenium thin films are typically grown directly on substrates and are thus more susceptible to substrate-induced stress, anisotropic grain growth, and structural disorder – all of which can significantly alter PL behavior. In 2022, Nielsen et al. reported two distinct PL emission bands at cryogenic temperatures from a selenium thin film integrated into a record-performing photovoltaic device [27]. Liu et al. observed similar spectral features, although the acquisition conditions were not specified [45]. Other studies have reported redshifted PL peak energies well below the optical bandgap in films infiltrated into mesoporous oxide scaffolds [26, 46]. Li et al. presented PL spectra with peak energies more reasonable for the optical bandgap, but without a detailed analysis of the nature of the underlying transition [1]. This inconsistency across the literature underscores the need for a more systematic investigation. However, the high vapor pressure of selenium makes thin films prone to sublimation and degradation during PL measurements, even under low excitation conditions. This challenge may have limited the widespread use of PL as a diagnostic tool, despite its value for evaluating optoelectronic quality.
In this work, we use temperature-dependent PL and Raman spectroscopy to study the optoelectronic and vibrational properties of selenium thin films synthesized in different laboratories using state-of-the-art fabrication methodologies. To address the challenge posed by the high vapor pressure and volatility of selenium, we adopt a recently developed closed-space encapsulation strategy introduced by Nielsen et al. [47], which effectively suppresses sublimation during the measurements. The vibrational properties are studied through a detailed analysis of the thermal evolution of Raman peaks, supported by first-principles phonon dispersion calculations that enable identification of all Raman-active modes and their associated atomic displacements. Temperature-dependent PL measurements, complemented by power-dependent excitation studies, are used to examine the nature of two emission bands and quantify the electron-phonon interactions that drive the thermal shifts and broadening of the observed radiative features. By comparing samples produced in different labs, we show that the short-range electron-phonon coupling strength is highly sensitive to subtle processing variations. This demonstrates that a reduction of microstructural disorder through careful control of growth parameters and thermal processing can substantially improve the optoelectronic quality of selenium thin films.
@@ II Results
Building on this group-theoretical framework, we next examine the phonon dispersion using first-principles density functional theory (DFT) calculations. Figure 2 shows the phonon dispersion and the total phonon density of states (DOS), with vibrational modes at the -point labeled by the symmetries introduced above. The A and doubly degenerate E(2) modes appear nearly degenerate in frequency at , suggesting they may be difficult to distinguish experimentally. The spectrum also reveals a pronounced phonon bandgap separating low- and high-frequency optical branches. While such a gap restricts available phonon scattering pathways – potentially suppressing phonon-phonon interactions and enhancing radiative efficiency – it may also channel energy from photoexcited carriers into specific high-energy modes, thereby increasing electron-phonon coupling and promoting non-radiative recombination. To modify or eliminate the phonon gap and thereby tailor intrinsic vibrational properties, phonon dispersion engineering through alloying, nanostructuring, or deliberate introduction of strain offers a promising strategy for enhancing optoelectronic performance potential.
Next, we deconvolute the experimental Raman spectrum acquired at room temperature using a rigorous methodology as detailed in Refs [51, 52], shown in Figure 3(a). The low-frequency peak at 146 cm corresponds well with the doubly degenerate E(1) mode predicted by the phonon dispersion. The peak at 235 cm overlaps with both the predicted E(2) and A modes, which are nearly degenerate and indistinguishable in energy. The slight asymmetry of this peak, particularly evident in the power-dependent Raman spectra in Figure 3(b), suggests contributions from two distinct but overlapping vibrational modes. However, due to the finite linewidth, limited spectral resolution, and low signal intensity, the two modes are best modeled using a single Lorentzian. These assignments are consistent with prior observations by Geick et al. [53]. Additionally, a broad multi-phonon band is observed between 430 and 475 cm, consistent with multi-phonon scattering involving 2A, 2E(2) and A + E(2) modes. The DFT-calculated phonon displacements of the three Raman-active modes are shown as an inset in Figure 3(a), demonstrating that the E(1) mode is a bending type vibration that shears the covalent bonds, whereas both E(2) and A correspond to asymmetric and symmetric bond-stretching vibrations, respectively. To better visualize these displacements, additional representations along different lattice directions are provided in Figure S1 of the Supplementary Information.
@@
As the temperature decreases, the E(1) mode is observed to blueshift. This is a vibrational fingerprint of interchain contraction, as the restoring force associated with the bond-shearing mode is strengthening, indicating an increased interchain coupling and a reduction in the distance between the selenium chains. In contrast, the E(2)+A peak, arising from two bond-stretching modes, redshifts upon cooling – opposite to what would be expected from simple lattice contraction. Although this redshift is smaller than the instrumental resolution, it is systematic and reproducible. This unexpected softening may result from the relaxation of grown-in compressive strain or the development of tensile strain along the chains at lower temperatures, both of which can reduce the effective restoring forces despite thermal contraction of the Se-Se bonds. Notably, this interpretation aligns with the temperature-dependent XRD measurements by Grosse et al. [54].
@@ III Discussion
The interplay between vibrational and optoelectronic properties has a strong influence on overall device performance. Our results show that structural disorder in selenium thin films is not intrinsic but highly sensitive to subtle variations in processing. These extended defects likely act as killer non-radiative recombination centers. Because such defects are not necessarily radiative, their impact may not be fully captured by PL measurements alone. However, we do observe a defect-bound exciton emission feature. Prior work by Nielsen et al. [59] identified an acceptor-like trap with a similar energy level, possibly corresponding to the (+1/0) charge transition level of selenium vacancies [28]. While this lower-energy exciton-related emission is unlikely to represent the primary non-radiative loss mechanism responsible for the large open-circuit voltage deficit, it may still contribute to the broad absorption onset commonly observed in both EQE and absorption spectra. This would help explain the persistent discrepancy between our PL-inferred optical bandgaps (1.84–1.86 eV), and the photovoltaic bandgaps (1.95 eV) derived from the inflection point of the EQE [49, 60, 61, 34]. This discrepancy is likely a result of sub-bandgap absorption from band tailing induced by traps and/or structural disorder. Earlier work by Chen et al. [62] also reported strong trapping of free excitons in selenium and proposed that such trapping could account for significant non-radiative recombination losses at higher temperatures. All of these observations are broadly consistent with the emerging picture of the microstructural and optoelectronic properties of selenium thin films.
@@ Acknowledgements
The work presented here is supported by the Carlsberg Foundation, grant CF24-0200. O.S.B. and E.S. acknowledge the CURIO-CITY project (PID2023-148976OB-C41); O.S.B., A.T. and M.P. acknowledge the INNO-PV project (PID2022-140226OB-C31C32); M.P. acknowledges the ENPOWER project (PCI2024-155100-2) from the CETP-Partnership Program 2023; all of them funded by MCIN/AEI/10.13039/501100011033/ FEDER. A.T. and M.P. acknowledge the SELECTRON project (CNS2023-14817) funded by MCIN/AEI/10.13039/501100011033/ NextGenerationEU/PRTR. Through our membership of the UK’s HEC Materials ChemistryConsortium, which is funded by the UK Engineering and Physical Sciences Research Council (EPSRC; EP/L000202, EP/R029431, EP/T022213), this work used the ARCHER2 UK National Supercomputing Services. We are also grateful to the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1). S.R.K thanks the Harvard University Center for the Environment for funding a fellowship. ES acknowledge the ICREA Academia Program.